OpenStudy (anonymous):

Find the equation of the line (2,-7) and (2,3). Thankssss

OpenStudy (anonymous):

first find the slope \[\Large m=\frac{3-(-7)}{2-2}=undefined\] so its vertical line !! can you find now the equation. ?? the equation of line passing through points (a,b) ,(a,c) is x=a can you now find it ?

OpenStudy (anonymous):

@isabelsantos can you now find it ?

hero (hero):

You can do that or you can just simply use y = mx + b. Would you like to know how @isabelsantos ?

OpenStudy (anonymous):

I got it, thank you so much!! :)

OpenStudy (anonymous):

yw:)

hero (hero):

You got it? What do you mean you got it? Mind explaining exactly what it is you got?

OpenStudy (anonymous):

I meant that I know how to do it :)

hero (hero):

I suppose, but I guess what I'm trying to say is, Can you prove to me that you know the steps? Perhaps, maybe by demonstrating them right here?

hero (hero):

That's the only way that I would be convinced that you truly know the steps.